OpenStudy (anonymous):

Is x+1 a factor of x^30 + 2x^17 +!? Explain. Please help! Is x+1 a factor of x^30 + 2x^17 +!? Explain. Please help! @Mathematics

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

i can't really understand what that thing is. is it \[x^{30}+2x^{17}+1\]?

OpenStudy (anonymous):

if so, then replace x by -1 and get \[(-1)^{30}+2\times (-1)^{17}+1=1-2+1=0\]

OpenStudy (anonymous):

since -1 gets you 0, the "factor theorem" tells you that x + 1 is a factor, so answer is yes.

OpenStudy (anonymous):

yes it is. Oh okay that makes sense thank you!

OpenStudy (anonymous):

yw